# Curious QNEIs from QNEC: New Bounds on Null Energy in Quantum Field Theory

### v1

@@ Acknowledgements.
$\epsilon$ is a coordinate width, and so becomes $\mathcal{A}/\sqrt{h}$ for higher dimensional curved manifolds, where $\mathcal{A}$ is the proper area of the top of the infinitesimal box shape variation. This then connects to the notation used in Bousso:2015mna .
@@ Appendix B Testing bounds with Bañados states
We will test some of our inequalities against states in holographic 2d CFTs dual to Bañados geometries Sheikh-Jabbari:2016znt . In these states, both the null energy density and the entanglement entropy are known.
@@
As mentioned in Sec. 2, all Bañados states saturate the 2d QNEC Khandker:2018xls ; Ecker:2019ocp ; what we have done here is verify that for constant $L_{v}$.

### v2

@@ Acknowledgements.
$\epsilon$ is a coordinate width, and so becomes $\mathcal{A}/\sqrt{h}$ for higher-dimensional curved manifolds, where $\mathcal{A}$ is the proper area of the top of the infinitesimal box shape variation. This then connects to the notation used in Bousso et al. (2016b).
@@ Appendix B Testing bounds with Bañados states
We will test some of our inequalities against states in holographic 2d CFTs dual to Bañados geometries Sheikh-Jabbari and Yavartanoo (2016). In these states, both the null energy density and the entanglement entropy are known.
@@
As mentioned in Sec. 2, all Bañados states saturate the 2d QNEC Khandker et al. (2018); Ecker et al. (2019); what we have done here is verify that for constant $L_{v}$.